OpenStudy (anonymous):

Jill wants to make a batch of ice tea in a cylindrical pitcher. The pitcher has a radius of 4.1cm and a height of 12.45cm. If the ice tea mixture only needs about 3/4 of water to fill the pitcher, how much water does Jill needs? Thank you

OpenStudy (anonymous):

The volume of a cylinder is \[\pi*r ^{2}* h\] So pi * (4.1)^2 * 12.45 = 657.49 cubic centimeters. Now you need 3/4 of that, so 657.49 * 3/4 = 493.11 Now convert 493.11 into whatever unit of water you're using.

OpenStudy (anonymous):

1 cm^3 = 1 milliliter